A box with a square base and no top is to be made from a square piece of cardboard by cutting 4 inch squares from each corner and folding up the sides. The box is to hold 100 in cubed. How big a piece of cardboard is needed.
Answer by cleomenius(959)

The volume will be L*W*H
We are cutting 4 inches from each corner, so the length and width will each be
x - 8.
The height will be 4.
100 = 4* (x - 8)* (x - 8)
25 = (x - 8)^2
5 = x - 8
13 inches= x.
The piece of cardboard needed will be 13 by 13 inches.
Cleomenius.